An engine 200 mm bore and 300 mm stroke works on Otto cycle. The clearance volume is 0.0016 m3. The initial pressure and temperature are 1 bar and 60°C. If the maximum pressure is limited to 24 bar, find : (i) The air-standard efficiency of the cycle (ii) The mean effective pressure for the cycle. Assume ideal conditions


1
Expert's answer
2021-07-16T03:56:13-0400

i.

"η=1-\frac{1}{r^{\gamma-1}} = 1-\frac{1}{8.85^{1.4-1}}=0.5879\\
\implies η= 58.19 \%"

ii.

"pm= \frac{P_1 r[r^{\gamma-1}-1]*[r_p-1]}{(r-1)(\gamma-1)}\\
pm= \frac{8.85[8.85^{1.4-1}-1]*[0.134]}{(0.4)(7.85)}\\
pm=0.526"